Hello I am stuck on this it doesnt appear to come oout any help would be appreciated

Simplify: (x^-7 y^11 z^14)^-2
John

Found 2 solutions by user_dude2008, checkley77:
Answer by user_dude2008(1862)

(x^-7 y^11 z^14)^-2 = x^(-7*-2)y^(11*-2)z^(14*-2) = x^14 y^(-22) z^(-28)

Answer by checkley77(12844)

(x^-7 y^11 z^14)^-2
X^14Y^-22Z^-28
X^14/Y^14Z^28